Graph : y=5(2)^x
thanks! show work please if some.
Last time i submitted it wrong, thanks for the help!

Answer by jim_thompson5910(35256)

To graph any equation, you need a set of points to plot. The more points you have, the more accurate the graph is.

To plot a point, we need the x and y coordinates of it. So if x = 0 for instance, then y is...

y = 5(2)^x
y = 5(2)^0 ... replace x with 0
y = 5*1
y = 5

If x = 0, then y = 5 leading to the point (0,5). This point lies on the curve.

Now if x = 1, then...

y = 5(2)^x
y = 5(2)^1 ... replace x with 1
y = 5*2
y = 10

If x = 1, then y = 10 leading to the point (1,10). This point lies on the curve as well.


Do this for other various values of x (x = 2, 3, 4, etc etc) to get corresponding paired values of y.

You will get this table

x y
0 5
1 10
2 20
3 40
4 80
5 160


To read this table, you just pick a row and read off the values in it in the form (x,y). Example: Pick on row 5. That has x = 5 and y = 160. So the point (5,160) lies on the curve.

Once you have enough points (the more points, the more accurate the curve is drawn), you draw a curve through them all to complete the graph.